# Object-Centric Image to Video Generation with Language Guidance

### v1

@@ Acknowledgment
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Limitations and Future Work
Prediction Artifacts   TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details in the objects, or visual artifacts in the background. We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the latent space representation.
Poor Temporal Consistency   We observe that TextOCVP’s predictions often lack temporal consistency when forecasting for long prediction horizons ($T>30$). We attribute this limitation to the fact that TextOCVP is trained by optimizing reconstruction losses, which do not penalize temporal inconsistencies.
Future Work   To address these limitations, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh et al., 2022) or diffusion models (Jiang et al., 2023), as well as scale our predictor module. Furthermore, we plan to incorporate temporal discriminators (Clark et al., 2020) to improve the temporal consistency of the predicted video frames. We believe that exploring these architectural modifications will enable us to utilize TextOCVP as a world model in complex real-world robotic environments.
@@ B.1 $\text{TextOCVP}_{\text{DINO}}$
$\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al., 2023) framework for recursive object-centric video decomposition and video rendering.
@@
Text-Conditioned Predictor   The predictor is composed of $N_{\text{Pred}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function. Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
@@
Text Encoder   $\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al., 2020), which consists of 6 T5 blocks. This text encoder uses a vocabulary with size 32,128.
@@
Scene Parsing   The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128. As feature extractor, we use DINOv2 ViT-Base (Oquab et al., 2023), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$. The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation. The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance. We use a single Transformer encoder block as the transition function, which consists of a four attention heads and an MLP with hidden dimension 512.
@@
Video Rendering   The video rendering module consists of two distinct decoders. First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al., 2019) with hidden dimension 1024 reconstructs the patch features from the slots. Then, to reconstruct the images from the features, we implement a CNN-based decoder with four convolutional layers, where each layer uses $3\times 3$ kernels. Every layer is followed by bilinear upsampling and a ReLU activation function. A final convolutional layer is applied to map to the RGB channels of the image.
@@
Training   We train our model for object-centric decomposition using videos sequences of length five frames for 1000 epochs. We use batch size of 16, the Adam optimizer (Kingma & Ba, 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given a frozen and pretrained object-centric decomposition model for 700 epochs. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
@@ B.2 $\text{TextOCVP}_{\text{SAVi}}$
$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al., 2022) as the object-centric decomposition module.
Scene Parsing   The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128. Following Kipf et al. (2022), we use as feature extractor a four-layer CNN with ReLU activation function, where each convolutional layer features 32 $5\times 5$ kernels, $\text{stride}=1$, and $\text{padding}=2$. The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
Video Rendering   Following Kipf et al. (2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al., 2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$, and a final convolutional layer which maps to four channels (RGB + alpha mask).
Training   We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given a frozen and pretrained object-centric decomposition model for 1400 epochs. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
@@ Appendix C Baselines
We employ four different baselines to compare againts our TextOCVP model on the image-to-video generation task on CATER and CLIPort datasets. To emphasize the importance of incorporating textual information, we include a comparison with OCVP-Seq (Villar-Corrales et al., 2023), a recent object-centric video prediction model that does not utilize text conditioning. Additionally, we evaluate a non-object-centric TextOCVP variant (Non-OC TF) that processes the input image into a single high-dimensional slot representation, instead of multiple object-centric slots, thus allowing us to evaluate the effect of object-centric representations. Moreover, we compare TextOCVP with two popular text-conditioned image-to-video generation baselines that do not incorporate object-centricity: MAGE (Hu et al., 2022) and SEER (Gu et al., 2024). We train both models on CATER and CLIPort closely following the original implementation details [^1] https://github.com/Youncy-Hu/MAGE [^2] https://github.com/seervideodiffusion/SeerVideoLDM/tree/main.
C.1 MAGE
MAGE is an autoregressive text-guided image-to-video generation framework that utilizes a VQ-VAE (Russakovsky et al., 2015) encoder-decoder architecture to learn efficient visual token representations. A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation. For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$. Moreover, on CLIPort we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our $\text{TextOCVP}_{\text{DINO}}$ approach. This adjustment ensures a fair comparison and significantly enhances MAGE’s performance on CLIPort. We refer to this modified version as $\text{MAGE}_{\text{DINO}}$. Additionally, to align with our experimental setup, we omit the speed parameter.
@@ C.2 SEER
SEER is a diffusion-based model for language-guided video prediction. It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al., 2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics. For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al., 2021) into frame-specific sub-instructions. These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers. We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al., 2017), and further fine-tune it for a few epochs. We observed that incorporating a text loss enhanced SEER’s performance, while other hyper parameters were kept consistent with its original implementation.
@@ Appendix D Datasets
CATER   CATER (Girdhar & Ramanan, 2020) is a dataset that consists of long video sequences, each described by a textual caption. The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates. The text instruction describes the movement of specific objects through four atomic actions: ‘rotate’, ‘pick-place’, ‘slide’, and ‘contain’. The caption follows a template consisting of the subject, action, and an optional object or end-point coordinate, depending on the action. The movement of the objects starts at the same time step. Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
@@
In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs. It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined toruses. Furthermore, every object is described by its size (small, medium, or large), material (metal or rubber), and color (red, blue, green, yellow, gray, brown, purple, cyan, or gold if the object is the snitch), and this description is included in the textual caption. Every atomic action is available. The ‘rotate’ action is afforded by cubes, cylinders and the snitch, the ‘contain’ action is only afforded by the cones, while the other two actions are afforded by every object. Every video has between 3 and 8 objects, and two actions happen to different objects at the same time. The vocabulary size is 50.
@@
CLIPort   CLIPort (Shridhar et al., 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e. long videos whose motion is described by a textual video caption. There are many variants of the CLIPort dataset, but we focus on the Put-Block-In-Bowl variant. We generate 21000 video-caption pairs with resolution $336\times 336$. Every video contains 6 objects on a 2D table plane, and a robot arm. Objects can be either a block or a bowl, and there is at least one of them in every sequence. The starting position of each object is random, with the only constrain to be placed on the table. Each video describes the action of the robot arm picking a block, and putting it in a specific bowl. The video caption follows the template ‘put the [color] block in the [color] bowl’. Each individual object in the scene has a different color. In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the test set they can have blue, green, red, pink, purple, white, or orange color. The other 4 objects, called distractors, can have any color. During a video sequence, it can be possible that the robot arm goes out of frame, and comes back in later frames, thus requiring the model to leverage long range dependencies. The vocabulary size is 15.
Appendix E Additional Results

### v2

@@ A.1 Limitations
Prediction Artifacts
TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details, or visual artifacts in the background. We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the object-centric latent space representation.
Limited Temporal Consistency
While TextOCVP produces plausible predictions, we observe that its temporal consistency can degrade when forecasting for long prediction horizons ($T>30$), occasionally resulting in object jittering or instability. We attribute this limitation to the fact that TextOCVP is trained to predict only up to nine future frames ($T=9$) and only optimizing reconstruction losses, which do not penalize such temporal inconsistencies.
A.2 Future Work
In future work we aim to address these limitations. Our TextOCVP model is designed with a modular architecture, enabling for both improvements and flexible swapping of components—such as the parsing, predictor, or video rendering modules—to seamlessly improve the entire pipeline.
To address the prediction artifacts, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh, Wu, and Ahn 2022) or diffusion models (Jiang et al. 2023), as well as scale our predictor module. Furthermore, we plan to incorporate temporal discriminators (Clark, Donahue, and Simonyan 2020) to improve the temporal consistency of the predicted video frames.
We believe that exploring these architectural modifications can overcome the aforementioned limitations and will enable the scaling of TextOCVP to real-world robotic environments.
@@ B.1 $\text{TextOCVP}_{\text{DINO}}$
The $\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al. 2023) framework for recursive object-centric video decomposition and video rendering.
Text-Conditioned Predictor
@@
The predictor is composed of $N_{\text{P}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function. Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
Text Encoder
@@
$\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al. 2020), which consists of six T5 blocks. This text encoder uses a vocabulary with size 32,128.
Scene Parsing
@@
The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128. As feature extractor, we use DINOv2 ViT-Base (Oquab et al. 2024), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$. The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation. The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance. We use a single Transformer encoder block as the transition function, which consists of four attention heads and an MLP with a hidden dimension of 512.
Video Rendering
@@
The video rendering module consists of two distinct decoders. First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al. 2019) with hidden dimension 1024 reconstructs the patch features from the slots. Then, a CNN-based decoder reconstructs full-resolution images from these features. It consists of four convolutional layers, each using $3\times 3$ kernels, a ReLU activation function and bilinear upsampling. A final convolutional layer and bilinear interpolation are applied to map the outputs to the RGB channels and spatial dimensions of the image.
Training
@@
We train our model for object-centric decomposition using video sequences of length five frames for 1000 epochs. We use batch size of 16, the Adam optimizer (Kingma and Ba 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given the frozen and pretrained object-centric decomposition model for 700 epochs to predict the subsequent nine frames using a single seed frame. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
@@ B.2 $\text{TextOCVP}_{\text{SAVi}}$
$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al. 2022) as the object-centric decomposition module.
Scene Parsing
The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128. Following Kipf et al. (2022), we use as feature extractor a four-layer CNN with ReLU activation function, where each convolutional layer features 32 $5\times 5$ kernels, $\text{stride}=1$, and $\text{padding}=2$. The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
Video Rendering
Following Kipf et al. (2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al. 2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$. A final convolutional layer maps from the hidden 32-channel representation to four output channels (RGB + alpha mask).
Training
We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given the frozen and pretrained object-centric decomposition model for 1400 epochs to predict the subsequent nine frames using a single seed frame. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
Table 4: Number of learnable parameters in TextOCVP and baselines for experiments on the CLIPort dataset.
Model # Parameters
TextOCVP 33.76M
Non-OC 34.16M
$\text{MAGE}_{\text{DINO}}$ 32.11M
SEER 405.89M
@@ Appendix C Baselines
[…]
Table 5: Comparison of MAGE variants on CLIPort. $\text{MAGE}_{\text{DINO}}$ clearly outperforms the original MAGE variant.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
MAGE 7.116 0.453 0.713
$\text{MAGE}_{\text{DINO}}$ 23.723 0.940 0.064
[…]
MAGE is an autoregressive text-guided video prediction framework that utilizes a VQ-VAE (Van Den Oord, Vinyals et al. 2017) encoder-decoder architecture to learn efficient visual token representations. A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation. For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$.
To ensure a fair comparison with TextOCVP on CLIPort, we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our TextOCVP model. We refer to this modified version as $\text{MAGE}_{\text{DINO}}$. Table 5 presents a comparison between the original MAGE model and $\text{MAGE}_{\text{DINO}}$ on CLIPort. The results demonstrate that $\text{MAGE}_{\text{DINO}}$ significantly outperforms the original variant, enabling a fair comparison with TextOCVP and other baselines on this benchmark.
@@ C.2 SEER
SEER is a diffusion-based model for language-guided video prediction. It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al. 2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics. For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al. 2021) into frame-specific sub-instructions. These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers. We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al. 2017), and further fine-tune it for a few epochs. We observed that incorporating a text loss enhanced SEER’s performance, while other hyper-parameters were kept consistent with its original implementation.
@@ D.1 CATER
CATER (Girdhar and Ramanan 2020) is a dataset that consists of long video sequences, each described by a textual caption. The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates. The text instruction describes the movement of specific objects through four atomic actions: ‘rotate’, ‘pick-place’, ‘slide’, and ‘contain’. The caption follows a template consisting of the subject, action, and an optional object or end-point coordinate, depending on the action. The movement of the objects starts at the same time step. Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
@@
In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs, with video frames resized to $64\times 64$. It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined toruses. Furthermore, every object is described by its size (small, medium, or large), material (metal or rubber), and color (red, blue, green, yellow, gray, brown, purple, cyan, or gold if the object is the snitch), and this description is included in the textual caption. Every atomic action is available. The ‘rotate’ action is afforded by cubes, cylinders and the snitch, the ‘contain’ action is only afforded by the cones, while the other two actions are afforded by every object. Every video has between 3 and 8 objects, and two actions happen to different objects at the same time. The vocabulary size is 50.
D.2 CLIPort
@@
CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e. long videos whose motion is described by a textual video caption. There are many variants of the CLIPort dataset, but we focus on the Put-Block-In-Bowl variant. We generate 21000 video-caption pairs with resolution $336\times 336$. Every video contains 6 objects on a 2D table plane, and a robot arm. Objects can be either a block or a bowl, and there is at least one of them in every sequence. The starting position of each object is random, with the only constraint being that it must be placed on the table. Each video describes the action of the robot arm picking a block, and putting it in a specific bowl. The video caption follows the template ‘put the [color] block in the [color] bowl’. Each individual object in the scene has a different color. In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the evaluation set with unseen colors they can have blue, green, red, pink, purple, white, or orange color. The other 4 objects, called distractors, can have any color. During a video sequence, it can be possible that the robot arm goes out of frame, and comes back in later frames, thus requiring the model to leverage long range dependencies. The vocabulary size is 15.
@@
CATER1→9 CATER1→19 CLIPort1→9 CLIPort1→19
[…]
OCVP 29.08 0.874 0.078 4.16 28.11 0.854 0.101 8.08
Non-OC 29.68 0.874 0.092 3.04 28.39 0.849 0.112 8.62 23.44 0.901 0.184 8.13 20.14 0.872 0.210 13.23
SEER 22.05 0.723 0.245 11.23 16.05 0.535 0.299 17.29 21.01 0.887 0.141 6.80 11.30 0.622 0.331 8.29
MAGE 34.91 0.877 0.108 3.46 34.76 0.871 0.111 5.88 23.72 0.940 0.064 2.11 22.27 0.931 0.075 2.59
TextOCVP 32.98 0.922 0.036 2.16 31.29 0.902 0.044 5.09 26.99 0.950 0.062 1.36 23.88 0.931 0.078 2.23
Appendix E Evaluation Metrics
To measure TextOCVP’s video prediction performance and compare it with existing approaches, we evaluate the visual quality of predicted video frames using popular image- and video-based metrics.
PSNR and SSIM (Wang et al. 2004) measure pixel-wise and statistical differences between the predicted and ground-truth video frames, respectively.
LPIPS (Zhang et al. 2018) is a perceptual metric that measures the visual similarity between two images based on deep features from pretrained neural networks, usually VGG (Simonyan and Zisserman 2015). Unlike pixel-wise metrics, LPIPS compares activations from multiple layers and captures differences in texture, structure, and semantics, thus aligning closely with human perception.
While these metrics focus on frame-level visual quality, video prediction evaluation requires quantifying temporal consistency and motion realism. JEDi (Luo et al. 2025) measures the quality and realism of generated videos by comparing feature distributions of generated and real videos, capturing both visual fidelity and motion dynamics. We prefer JEDi over FVD (Unterthiner et al. 2019), as JEDi is less sensitive to small evaluation datasets, and more stable in low-motion synthetic scenarios.
In our evaluations and comparisons with baselines, we favor the LPIPS and JEDi scores, which correlate well with human perception, while reporting other metrics for completeness.
Appendix F Additional Results
F.1 SAVi vs. DINOSAUR
Current object-centric approaches for video prediction are limited to relatively simple synthetic datasets, and struggle to scale beyond scenes featuring simple 3D shapes with simple deterministic motion. We attribute this limitation primarily to the object-centric modules used for learning object representations. Motivated by this observation, we extend the recent DINOSAUR (Seitzer et al. 2023) framework to recursively process video data and reconstruct video frames from their corresponding object-centric representations.
To demonstrate the significance of the object-centric module in scaling to more complex datasets, we compare both SAVi and our Extended DINOSAUR trained on CLIPort. As illustrated in Fig. 8, SAVi struggles to accurately represent the objects on the table, missing multiple objects and changing their shape and color. In contrast, the Extended DINOSAUR model successfully reconstructs the scene, closely resembling the input, while accurately representing each object. The visual features extracted by the DINOv2 (Oquab et al. 2024) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information. This design enables the model to scale and handle more complex object-centric video data effectively.
$t=1$
22
33
44
55
GT
SAVi
Ext. DINO